OpenStudy (anonymous):

a 35 foot telephone pole casts a 46 foot shadow on the ground while the sun is shining. to the nearest degree, what is the angle of elevation of the sun from the end of the shadow.

OpenStudy (anonymous):

well which trig identity solves for a x and a y direction. That would be tangent right. Solve for your value by moving it over and doing tan^-1 to get your angle.

OpenStudy (anonymous):

Answer @megss_69: tanX= 35/46 x=tan-1 (35/46) X=37 degrees